OpenStudy (anonymous):

Which equation is related to the equation shown? x + 3 = 12 A. x = 12 + 3 B. x = 12 ÷ 3 C. x = 12 − 3 D. x = 12 • 3

OpenStudy (anonymous):

can someone please help i will fan and medal

OpenStudy (anonymous):

It will be C

OpenStudy (anonymous):

because it will be backword

OpenStudy (anonymous):

i thought so thanks

OpenStudy (anonymous):

can you help with another one

OpenStudy (anonymous):

Which is an equation for the given word problem? After Elise deposits $15.50 into her bank account, her balance is $325.00. Which equation can be used to find the amount, n, in her account before she made the deposit? A. 15.50 – n = 325.00 B. 15.50 + n = 325.00 C. n + 325.00 = 15.50 D. n – 325.00 = 15.50

OpenStudy (anonymous):

i beleive it is b but i am not sure

OpenStudy (anonymous):

hello